OpenStudy (gabylovesyou):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $130. This year, he paid $143. What was the percent increase in the cost of his plane ticket? Show your work.

jimthompson5910 (jim_thompson5910):

Percent Increase = [(New - Old)/Old] * 100 Percent Increase = [(143-130)/130] * 100 Percent Increase = [13/130] * 100 Percent Increase = 0.1 * 100 Percent Increase = 10% So the percent increase of the cost of his plane ticket was 10% Note: this means that you add 10% of 130 onto 130 to get the new price of $143

OpenStudy (gabylovesyou):

is it 10% increase?

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (gabylovesyou):

lol just making sure my answer was correct!! :D lol

jimthompson5910 (jim_thompson5910):

you're on a roll lol

OpenStudy (gabylovesyou):

lol thank you!!

jimthompson5910 (jim_thompson5910):

lol yw

OpenStudy (gabylovesyou):

u get lots of medals from me! lol everytime u want a medal u come lol

jimthompson5910 (jim_thompson5910):

lol thx for the offer, I'll be sure to keep that in mind

OpenStudy (gabylovesyou):

ok :D

OpenStudy (anonymous):

10% because difference b/w two tickets is 13 and by dividing this to 130 we got the percentage increased. ok

OpenStudy (gabylovesyou):

thank you! @nitish1992